OpenStudy (anonymous):

A gas has a pressure of 7.00 atm at 227 C. What will its pressure be at 27 C if the volume does not change?

OpenStudy (anonymous):

\[T _{1}=227+273=500~k\] \[T _{2}=27+273=300~K\] \[\frac{ P _{1} }{ T _{1} }=\frac{ P _{2} }{ T _{2} }\]

OpenStudy (anonymous):

What does T represent?

OpenStudy (anonymous):

Temperature, sorry

OpenStudy (anonymous):

temperature in Kelvin scale

OpenStudy (anonymous):

4.2 atm

OpenStudy (perl):

you actually dont need to use kelvin here, but its a good habit to get into , because of the other equation PV = nRT , must use kelvin

OpenStudy (perl):

, when R is .08206